Cesar plans to retire at age 57, and estimates that $62,400 a year in retirement for 28 years will be needed. That annual amount will be paid out in monthly income checks. To save for retirement, Cesar plans to put money every month into a 401k plan at work. If Cesar is currently 25, how much will need to be saved each month to meet the retirement goal, assuming the retirement accounts earn an average of 6.6% per year.

Please show your work.
Answer #1

Amount needed a year in retirement= $62,400

Therefore, amount needed a month= $62,400/12 = $ 5,200

Amount needed at the time of retirement (age 57) is the present value of annuity as on that date, calculated at $795,740.18 as follows:

A B C D 1 Present Value of Annuity Payments at the end of each period 2 3 Present value of annuity is calculated using the fo

Period left till age 57 = 57-25= 22 years

Amount needed to be saved each month in order to accumulate this amount is $1,344.70 calculated as follows:

22 A B C D E 3 Amount of periodical payments is calculated using the formula PMT= (PV*r)/(1-(1+r)^-n] 4 Where PMT= Periodical